The potential energy of a particle in a certain field has the formU =a/r²-b/r,where  A and B are positive constants, r is the distance from the centre of the field. the value of r0 corresponding to the equilibrium position of the particle (state if it is stable or unstable )is

Grade:11

1 Answers

Arun
5 years ago
for equilibrium dU/ dr = 0
-2A/r³ + B/r² = 0 r = 2A / B
for stable equilibrium d²U /dr² should be positive for the value of r.
here d²U / dr² = 6A/r⁴ – 2B/r³ is +ve value for r = 2A / B So.